OpenStudy (anonymous):

Completely factor 9x^2-36a^4

OpenStudy (cwrw238):

note that 9x^2 and 36a^4 are perfect squares - we have the difference of 2 squares so we use the identity a^2 - b^2 = (a + b)(a - b) where a = 3x and b = 6a^2

OpenStudy (anonymous):

What's next?

OpenStudy (cwrw238):

thats (3x + 6a^2)(3x - 6a^2) now we can take out 3 = 3(x + 2a^2)* 3(x - 2a^2) = 9(x + 2a^2)(x - 2a^2)

OpenStudy (anonymous):

Thanks! Can you help me with a couple more?